OpenStudy (minecraftqueen):

i need help asap ill medal and fan if there right What is the value of n? 3n+6/3=15 2.The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle?

OpenStudy (anonymous):

n=4.33

OpenStudy (anonymous):

width of rectangle is 14 inches

OpenStudy (anonymous):

For the first problem if (3n+6)/3=15 then n=13 However, if 3n+(6/3)=15 then n=13/3

OpenStudy (anonymous):

For the second problem w=l-6 and l+w=68 Therefore, l+(l-6)=68. Thus, l=37 and w=37-6=31 Length=37 Width=31